OpenStudy (anonymous):

I'm taking the math SAT tomorrow, please help with this last question: "a Piece of Clay is shaped into a container with a volume of 36pi. If the cylinder is rolled into a sphere, what will the sphere's radius be? " The answer is 3, but how do you get that?

OpenStudy (dumbcow):

set volumes equal cylinder = pi*r^2*h sphere = 4/3 pi*r^3 4/3 pi*r^3 = 36pi pi cancels 4/3*r^3 = 36 r^3 = 36*(3/4) r^3 = 27 r = cuberoot(27) r = 3

OpenStudy (anonymous):

OMG, Thank You SO Much!!!! I can't believe I didn't see that. Thank You <3